OpenStudy (anonymous):

6x - 2y = 50 x + 3/2y = 12 Solve Using Substition @Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

alright :) so first, manipulate the 2nd equation so that it equals x

HanAkoSolo (jamierox4ev3r):

like so x+3/2y=12 -3/2y -3/2y \(x=\Large\frac{-3}{2}y\)

HanAkoSolo (jamierox4ev3r):

i mean ** \(x=\large\frac{-3}{-2}y + 12\)

HanAkoSolo (jamierox4ev3r):

so then we plug that in to the second equation \(6(\large\frac{-3}{2}+12)\)-2y=50

OpenStudy (anonymous):

Ok

HanAkoSolo (jamierox4ev3r):

so then distribute :) \(6(\large\frac{-3}{2}+12)-2y=50\) -9+72-2y=50

HanAkoSolo (jamierox4ev3r):

from here, add like terms -9+72-2y=50 63-2y=50

HanAkoSolo (jamierox4ev3r):

then subtract 63 from both sides 63-2y=50 -63 -63 -------- -2y=-13

HanAkoSolo (jamierox4ev3r):

lastly, divide -13 by -2y. then you'll be able to solve for y :)

HanAkoSolo (jamierox4ev3r):

@Ami22 I hope you've found this helpful! I will tell you how to solve for x in a little bit, alright? my mom wants me to go pick up a few things from the store. In the meantime, could you medal me? All you have to do is click on the blue best response button next to my name. Thank you :)

OpenStudy (anonymous):

ok